The monopoly firm has two consumers: Ali and Javad. Both have different demand equations. Ali's demand equation Q = 100-P; Javad's Demand equation Q = 120-P. The marginal cost of the firm is 20 manat. Using two-tier pricing, the firm has set a marginal cost per user and an additional 30 manat. Calculate the firm's profit and total profit for each consumer
1
Expert's answer
2020-01-08T09:34:49-0500

The profit is maximized, when such output is produced, at which MR = MC.

For Ali:

P = 100 - Q, MR = TR'(Q) = 100 - 2Q,

100 - 2Q = 50,

Q = 25 units, P = 100 - 25 = 75 manat.

TP = (75 - 50)×25 = 625 manat.

For Javad:

P = 120 - Q, MR = 120 - 2Q,

120 - 2Q = 50,

Q = 35 units, P = 120 - 35 = 85 manat.

TP = (85 - 50)×35 = 1225 manat.

Total profit is: TP = 1225 + 625 = 1850 manat.